OpenStudy (anonymous):

Amanda is taking a quiz that consists of five multiple choice questions, each having three answers from which to choose. If she guesses on each question, what's the probability that Amanda will get 100% on the quiz? 1/12 1/15 1/125 1/243

OpenStudy (anonymous):

1/15 right because 1/3 is the chance and there is five so 3 times five is 15

jimthompson5910 (jim_thompson5910):

There are 3^5 = 243 ways to answer 5 multiple choice questions (3 choices per question). There is only one way to get 100% So the probability is 1/243

OpenStudy (anonymous):

(1/3)(1/3)(1/3)(1/3)(1/3)=1/243

OpenStudy (anonymous):

thanks i see how u did that now i was just thinking about one of them

OpenStudy (anonymous):

No \[\frac{1}{3}.\frac{1}{3}.\frac{1}{3}.\frac{1}{3}.\frac{1}{3}\] there is one third chance of getting correct answers in each question so prbablities multiply as they are simultaneous and independent events